OpenStudy (anonymous):

This identity means that translating the basic sine graph Π/2 units to the right produces a cosine graph. sin(Π/2-x) = cosx

OpenStudy (anonymous):

true or false

OpenStudy (freckles):

Draw a right triangle and compare ratios

OpenStudy (freckles):

 that other angle is (pi/2-x) since (pi/2-x)+x must equal pi/2

OpenStudy (freckles):

now find cos(x) and sin(pi/2-x)

OpenStudy (freckles):

and i guess you could always use the difference identity for the sin(pi/2-x) if you don't like this triangle idea

ganeshie8 (ganeshie8):

i think the given statement is wrong translating sinx to right by pi/2 should give sin(x-pi/2)

OpenStudy (anonymous):

Remember that \(\cos = \text{co-sin}\). There the meaning for \(\text{co}\) that \(\text{co}f(x) = f(\pi/2-x)\)

OpenStudy (anonymous):

This is why \(\cos(x) = \sin(\pi/2-x)\). It's almost a definition.

OpenStudy (freckles):

didn't read the paragraph above about the identity

OpenStudy (freckles):

I agree with the guru @ganeshie8

OpenStudy (freckles):

now if it had said translating the basic graph of -sin graph pi/2 units to the right produces a cosine graph I might feel better about that

OpenStudy (freckles):

use the fact the sin is odd and that sin and cos are cofunctions

OpenStudy (freckles):

I think let me see

OpenStudy (freckles):

\[-\sin(x-\frac{\pi}{2}) \\ =\sin(-x+\frac{\pi}{2}) =\sin(\frac{\pi}{2}-x)\]

OpenStudy (anonymous):

so its a false statement?

ganeshie8 (ganeshie8):

yes i would tick false as it is not a direct consequence of that identity

ganeshie8 (ganeshie8):

sin(pi/2-x) = cos(x) can be thought of like this : reflection of sin(x) about y axis and then a translation of pi/2 units to the left